6. A prime p is called a Mersenne/Fermat prime if it has what form?
Answer #1

い! eimye
Answer #2

ANSWER :


Mersenne primes are of the form : 2^n - 1 ; n = natural number ≥ 1 .

(excluding 2^4 - 1 = 15 which is not a prime).


Fermat primes are of the form : 2^n + 1 ; n =  any natural number.

(excluding 2^3 + 1 = 9 which is not a prime)

Fermat’s more popular form : 2*2^n + 1 ; n = any natural number.

answered by: Tulsiram Garg